# A Survey on Trustworthiness in Foundation Models for Medical Image Analysis

### v1

@@ 4-E Disease Diagnosis
4-E1 Privacy in Disease Diagnosis
@@
Due to foundation models for medical image analysis typically trained on large-scale datasets which always contain sensitive patient private information, privacy considerations in foundation models for medical image analysis are increasingly paramount. As a gold standard framework for privacy-preserving, Differential privacy (DP) can provide a privacy guarantee for machine learning training without information leakage. However, degradations in model performance pose an obstacle to its application. To tackle this challenge, a privacy-preserving approach is proposed, involving the fine-tuning of pre-trained foundation models with differential privacy (DP) [70]. Numerous experimental results indicate that this approach can achieve comparable accuracy to non-private classifiers for medical image analysis, even in the presence of substantial distribution shifts between pre-training data and downstream tasks. In addition, privacy threats can also occur in the data sharing process before foundation model training. By introducing a stable data distillation method for medical image analysis through a progressive trajectory matching strategy, this approach offers a privacy-preserving data-sharing mechanism for pretraining foundation models [89]. Another solution for privacy leakage caused by data sharing is to medical artificial intelligence-generated content (Med-AIGC) as an inexhaustible resource repository. The utilization of artificially synthesized images along with corresponding textual data for training enables the medical foundation model to effectively absorb knowledge of disease symptomatology, thereby mitigating potential breaches of patient confidentiality [75]. To satisfy “the right to be forgotten” in the regulations [90], seamless data modification is utilized to improve privacy in medical image analysis without model retraining, thus reducing the risk of privacy breaches [73].
4-E2 Robustness in Disease Diagnosis
@@
Although foundation models for medical image analysis achieved great success, the robustness of foundation models for medical image analysis is still considerable. Veerla et al. [91] explores the vulnerabilities of the Pathology Language-Image Pretraining (PLIP) model by employing Projected Gradient Descent (PGD) adversarial attacks to intentionally induce misclassifications. The findings of the study emphasize the pressing need for robust defenses to ensure the security of foundation models for medical image analysis. Jin et al. [31] point out that backdoor attacks pose a threat to MedClip and identify vulnerabilities like BadMatch, which exploits minor label discrepancies. The study reveals that current defenses are inadequate against such backdoor attacks in medical foundation models. Another study also finds out that manipulating just 1.1% of an LLM’s weights allows for the injection of inaccurate biomedical facts that propagate throughout its output without impacting its performance on other tasks [92]. This susceptibility highlights serious security and trustworthiness concerns regarding the utilization of LLMs in healthcare.
4-E3 Reliability in Disease Diagnosis
Similarly, ensuring the reliability of foundation models in medical image analysis is crucial, necessitating measures to prevent hallucination and the propagation of outdated information. GPT-4 demonstrates significantly enhanced capabilities compared to its predecessors, particularly in the healthcare domain where its advanced abilities are prominently showcased. Wu et al. aim to assess the performance of GPT-4 by OpenAI in multimodal medical diagnosis, covering 17 human body systems (e.g., , Central Nervous System, Cardiac, Hematology, Oncology, Pediatrics) and 8 imaging modalities (e.g., X-ray, CT, MRI, PET, DSA, Mammography, Ultrasound, Pathology) [67]. Their evaluation focuses on various clinical tasks, including imaging modality recognition, disease diagnosis, report generation, and disease localization, with or without patient history.
@@ 4-E3 Reliability in Disease Diagnosis
In practical cases, GPT-4 shows its safety guarantees against potential misuse and ensures users are aware of its capabilities and limitations. Specifically, when GPT-4 is asked to make a diagnosis, such as providing a diagnosis for a chest X-ray, it incorporates safeguards: refusal to offer a diagnosis, emphasis on limitations, and expression of uncertainty.
4-E4 Explainability in Disease Diagnosis
@@
As a high-stakes domain with rigorous demands of trustworthiness, foundation models used for disease diagnosis must not only demonstrate high performance but also fulfill the criteria for explainability. To meet the explainable requirements and inspired by description-based interpretable approach [93], one recent study proposes a framework for explainable zero-shot medical image classification utilizing vision-language models like CLIP along with LLMs like ChatGPT [66]. The primary concept involves harnessing ChatGPT’s capabilities to automatically generate comprehensive textual descriptions encompassing disease symptoms and visual attributes, moving beyond mere disease labels. This supplementary textual data enhances the precision and interpretability of diagnoses generated by CLIP. By aligning the semantics of images, learnable prompts, and clinical concept-driven prompts at various levels of detail, Bie et al. [71] propose a novel explainable prompt learning framework, leveraging medical expertise. Agarwal et al. [72] presents a novel explainability strategy in healthcare, leveraging a vision-language model to identify language-based descriptors of visual classification tasks, demonstrating alignment with clinical knowledge and potential for aiding non-expert human understanding of specialized medical tasks. In pursuit of explainability in deep learning, Doerrich et al. [73] integrate a k-Nearest Neighbor (k-NN) classifier with a vision-based foundation model, enhancing both interpretability and adaptability. Yan et al. [74] propose a new paradigm to build interpretable medical image classifiers with natural language concepts, wherein they query clinical concepts from GPT-4 and transform latent image features into explicit concepts using a vision-language model.
4-E5 Fairness in Disease Diagnosis
@@
In the realm of disease diagnosis, ensuring fairness is significant as we delve into the intricate interplay between medical algorithms and demographic attributes. Some experimental results demonstrate that some foundation models for healthcare are considered unfair to some extent. Chen et al. [12] examine their proposed foundation model, CheXagent, for Chest X-ray interpretation, conducting a fairness evaluation, in which they examine different groups of people according to their sex, race and age. The results unveil disparities in model performance across various groups and underscore the presence of biases. Furthermore, Yang et al. [68] investigate the fairness of the existing Vision-language (VL) foundation models in chest X-ray diagnosis across five globally-sourced datasets. Their findings unveil a consistent pattern: these VL models consistently underdiagnose marginalized groups, with even higher rates observed in intersectional subgroups, such as Black female patients, when compared to diagnoses by board-certified radiologists.
@@
Inspired by some widely-used VL models such as CLIP, VL models for medical tasks such as MedCLIP also have attracted a lot of attention [2]. However, the absence of medical VL datasets poses a significant challenge for studying fairness. Luo et al. [69] addressed this issue by introducing Harvard-FairVLMed dataset. This dataset offers comprehensive demographic attributes, ground-truth labels, and clinical notes, enabling a thorough investigation into fairness within VL foundational models. Based on Harvard-FairVLMed dataset, some fairness issues are found when using two widely-used VL models (CLIP and BLIP2). To address these, an optimal-transport-based method named FairCLIP is proposed. It aims to trade off balance between model performance and fairness by minimizing the distribution gaps across both the overall dataset and individual demographic groups. Considering the scenario of using foundation model API, in which we have very limited model control and computational resources, Jin et al. [94] propose a Universal Debiased Editing (UDE) strategy. This method is applicable to both white-box and black-box foundation model APIs, with the capability to mitigate bias within both the foundation model API embedding and the images themselves. In addition, the fairness in privacy-preserving foundation models for medical images is investigated. Berrada et al. [70] examine pre-trained foundation models fine-tuned with Differential Privacy (DP) using two medical imaging benchmarks. Evaluation results show that private medical classifiers do not exhibit larger performance disparities across demographic groups than non-private models, making DP training a practical and reliable approach.
@@ 5-A Datasets and Benchmarks
5-B Hallucination
@@ 5-C Alignment
5-D Regulatory and Ethical Frameworks
@@ 6 Conclusion
In conclusion, our comprehensive exploration of foundation models within medical image analysis underscores the potential trustworthiness issues associated with applications and the use of foundation models. Foundation models have presented their transformative potential to significantly advance various medical imaging tasks, including segmentation, report generation, medical Q&A, and disease diagnosis. Despite these promising advancements, our investigation also highlights critical concerns regarding the trustworthiness of these foundation models for medical image analysis, particularly in terms of privacy, robustness, reliability, explainability and fairness. By reviewing the existing literature on foundation models for medical image analysis, we analyze the trustworthiness in diverse applications, pinpointing where these models excel and where they falter from trustworthy perspectives. This survey maps the current state of foundation models in medical image analysis and outlines critical challenges and directions for future research, emphasizing the need for trustworthy advancements. Addressing these issues is crucial for fully realizing the potential of foundation models in enhancing healthcare outcomes, fostering a future where innovation is coupled with ethical responsibility and equity.
Acknowledgements
This work is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), CIFAR Catalyst Grant, and CIFAR AI Chair awards.
Author Contributions
C.S., R.R., and J.Y. are co-first authors, each having conducted extensive portions of the literature review, drafted various sections of the manuscript, and collaborated on the thematic categorization and comparative analyses. Q.D., provides critical guidance, supervision, editing, and a comprehensive final review of the manuscript to ensure its scholarly quality. X.L. is the corresponding author, having conceived the idea for the survey, secured funding support, led the project, and contributed to writing and revising the manuscript. All authors have read and approved the final version of the paper.
Competing Interests
The authors declare no competing interests.

### v2

@@ 4.4 Disease Diagnosis
4.4.1 Privacy in Disease Diagnosis
@@
Trade-off problem when using differential privacy for model fine-tuning. Due to the fact that foundation models for medical image analysis are typically trained on large-scale datasets which may contain sensitive information from patients, privacy considerations in foundation models for medical image analysis are increasingly paramount. As a gold standard framework for preserving privacy, differential privacy (DP) can provide privacy guarantees for machine learning training without information leakage. However, degradations in model performance pose an obstacle to its application. To tackle this challenge, a privacy-preserving approach is proposed, involving the fine-tuning of pre-trained foundation models with differential privacy (DP) [71]. Numerous experimental results indicate that this approach can achieve comparable accuracy to non-private classifiers for medical image analysis, even in the presence of substantial distribution shifts between pre-training data and downstream tasks.
Data distillation and data synthesis for privacy preservation. In addition, privacy threats can also occur in the data sharing process before foundation model training. By introducing a stable data distillation method for medical image analysis through a progressive trajectory matching strategy, this approach offers a privacy-preserving data-sharing mechanism for pretraining foundation models [90]. Another solution for privacy leakage caused by data sharing is to use synthesized images. The utilization of artificially synthesized images along with corresponding textual data for training enables the medical foundation model to effectively absorb knowledge of disease symptomatology, thereby mitigating potential breaches of patient confidentiality [76]. To satisfy “the right to be forgotten” in the regulations [91], seamless data modification is utilized to improve privacy in medical image analysis without model retraining, thus reducing the risk of privacy breaches [74].
4.4.2 Robustness in Disease Diagnosis
@@
Vulnerabilities to adversarial attacks, backdoor attacks, and weight manipulations. Although foundation models for medical image analysis achieved great success, the robustness of foundation models for medical image analysis is still considerable. Veerla et al. [92] explores the vulnerabilities of the Pathology Language-Image Pretraining (PLIP) model by employing Projected Gradient Descent (PGD) adversarial attacks to intentionally induce misclassifications. The findings of the study emphasize the pressing need for robust defenses to ensure the security of foundation models for medical image analysis. Jin et al. [31] point out that backdoor attacks pose a threat to MedCLIP and identify vulnerabilities like BadMatch, which exploits minor label discrepancies. The study reveals that current defenses are inadequate against such backdoor attacks in medical foundation models. Another study also finds out that manipulating just 1.1% of an LLM’s weights allows for the injection of inaccurate biomedical facts that propagate throughout its output without impacting its performance on other tasks [93]. This susceptibility highlights serious security and trustworthiness concerns regarding the utilization of LLMs in healthcare.
4.4.3 Reliability in Disease Diagnosis
@@
Satety evaluation of applying GPT-4 for medical domain. Similarly, ensuring the reliability of foundation models in medical image analysis is crucial, necessitating measures to prevent hallucination and the propagation of outdated information. Wu et al. [68] assess the performance of GPT-4 by OpenAI in multimodal medical diagnosis, they found that GPT-4 shows its safety guarantees against potential misuse and ensures users are aware of its capabilities and limitations. Specifically, when GPT-4 is asked to make a diagnosis, such as providing a diagnosis for a chest X-ray, it incorporates safeguards: refusal to offer a diagnosis, emphasis on limitations, and expression of uncertainty.
4.4.4 Explainability in Disease Diagnosis
@@
Enhancing explainability by concept-bottleneck models. As a high-stakes domain with rigorous demands of trustworthiness, foundation models used for disease diagnosis must not only demonstrate high performance but also fulfill the criteria for explainability. To meet the explainable requirements and inspired by description-based interpretable approach [94], one recent study proposes a framework for explainable zero-shot medical image classification utilizing vision-language models like CLIP along with LLMs like ChatGPT [67]. The primary concept involves harnessing ChatGPT’s capabilities to automatically generate comprehensive textual descriptions encompassing disease symptoms and visual attributes, moving beyond mere disease labels. This supplementary textual data enhances the precision and interpretability of diagnoses generated by CLIP. By aligning the semantics of images, learnable prompts, and clinical concept-driven prompts at various levels of detail, Bie et al. [72] propose a novel explainable prompt learning framework, leveraging medical expertise. Agarwal et al. [73] presents a novel explainability strategy in healthcare, leveraging a vision-language model to identify language-based descriptors of visual classification tasks, demonstrating alignment with clinical knowledge and potential for aiding non-expert human understanding of specialized medical tasks. In pursuit of explainability in deep learning, Doerrich et al. [74] integrate a k-Nearest Neighbor (k-NN) classifier with a vision-based foundation model, enhancing both interpretability and adaptability. Yan et al. [75] propose a new paradigm to build interpretable medical image classifiers with natural language concepts, wherein they query clinical concepts from GPT-4 and transform latent image features into explicit concepts using a vision-language model. Wang et al. [95] introduce CopilotCAD, which combines the foundation model’s computational power with the expertise of radiologists, offering a user-friendly interface for interactive, image-based diagnostics. It empowers radiologists to make informed decisions with the support of AI-generated quantitative data and visual aids, enhancing the explainability and transparency of CAD systems.
4.4.5 Fairness in Disease Diagnosis
@@
Disparate model performance on marginalized groups. In the realm of disease diagnosis, ensuring fairness is significant as we delve into the intricate interplay between medical algorithms and demographic attributes. Some experimental results demonstrate that some foundation models for healthcare are considered unfair to some extent. Chen et al. [5] examine their proposed foundation model, CheXagent, for Chest X-ray interpretation, conducting a fairness evaluation, in which they examine different groups of people according to their sex, race and age. The results unveil disparities in model performance across various groups and underscore the presence of biases. Furthermore, Yang et al. [69] investigate the fairness of the existing Vision-language (VL) foundation models in chest X-ray diagnosis across five globally-sourced datasets. Their findings unveil a consistent pattern: these VL models consistently underdiagnose marginalized groups, with even higher rates observed in intersectional subgroups, such as Black female patients, when compared to diagnoses by board-certified radiologists.
@@
New dataset for fairness evaluation fairness Inspired by some widely-used VL models such as CLIP, VL models for medical tasks such as MedCLIP also have attracted a lot of attention [4]. However, the absence of medical VL datasets poses a significant challenge for studying fairness. Luo et al. [70] addressed this issue by introducing Harvard-FairVLMed dataset. This dataset offers comprehensive demographic attributes, ground-truth labels, and clinical notes, enabling a thorough investigation into fairness within VL foundational models. Based on Harvard-FairVLMed dataset, some fairness issues are found when using two widely-used VL models (CLIP and BLIP2).
To mitigate fairness issues, methods such as FairCLIP and Universal Debiased Editing are proposed. To improve the fairness of VL models, an optimal-transport-based method named FairCLIP is proposed. It aims to trade off balance between model performance and fairness by minimizing the distribution gaps across both the overall dataset and individual demographic groups. Considering the scenario of using foundation model API, in which we have very limited model control and computational resources, Jin et al. [96] propose a Universal Debiased Editing (UDE) strategy. This method is applicable to both white-box and black-box foundation model APIs, with the capability to mitigate bias within both the foundation model API embedding and the images themselves. In addition, the fairness in privacy-preserving foundation models for medical images is investigated. Berrada et al. [71] examine pre-trained foundation models fine-tuned with Differential Privacy (DP) using two medical imaging benchmarks. Evaluation results show that private medical classifiers do not exhibit larger performance disparities across demographic groups than non-private models, making DP training a practical and reliable approach.
@@ 5.1 Datasets and Benchmarks
Currently, FairMedFM [97] is proposed to address the lack of comprehensive benchmarks by offering an integrated framework that evaluates fairness in medical imaging foundation models. By integrating 17 popular medical imaging datasets and analyzing 20 widely used foundation models, it reveals the existence of bias, fairness-utility tradeoffs, inherent fairness issues in the datasets, and limited effectiveness of existing mitigation methods.
5.2 Hallucinations
@@ 5.3 Alignment
5.4 Regulatory and Ethical Frameworks
@@ 6 Conclusion
In conclusion, this work presents the first comprehensive analysis of foundation models in medical image analysis, focusing on the potential trustworthiness issues essential for responsible healthcare applications. Foundation models show transformative potential, enhancing tasks like segmentation, report generation, medical Q&A, and disease diagnosis. However, our survey highlights significant concerns regarding their trustworthiness, particularly in privacy, robustness, reliability, explainability, and fairness.
We categorize and evaluate trustworthiness across these dimensions, offering a nuanced view of where foundation models excel and where limitations persist. This analysis contributes a foundational map of the current state of foundation models in medical image analysis, outlining critical challenges and actionable directions for future research. Specifically, addressing these trustworthiness challenges is paramount to harnessing foundation models’ full potential to improve healthcare outcomes.
The path forward requires efforts on building standardized datasets and benchmark for evaluation, innovations that align technical progress with ethical responsibility and equity, ensuring that foundation models in healthcare not only advance performance but also adhere to principles that safeguard patient trust and societal impact.